OpenStudy (anonymous):

Solve 6cos^2x-sinx-4=0

OpenStudy (saifoo.khan):

6[cos²(x)] − sinx − 4 = 0 substitute [cos²(x)] = 1 − sin²(x) substitute u = sin(x) solve quadratic for "u" ... and so on

OpenStudy (anonymous):

After solving for u, make sure that it lies between -1 and 1, because any other value of u cannot be equal to sin(x) or cos(x). Discard any solution that isn't in the interval.

OpenStudy (anonymous):

But I thought cos2x=1-2sin^2x....is cox^2x=1-sin^2x as well?

OpenStudy (anonymous):

\[ \sin^2 x + \cos^2 x = 1 \] Rearrange to get cos²(x) in terms of sin²(x).

OpenStudy (anonymous):

oh right okay thank you:)